# Efficient Approximation Algorithms for Fair Influence Maximization under Maximin Constraint

### v1

@@ Abstract
Aiming to reduce disparities of influence across different groups, Fair Influence Maximization (FIM) has recently garnered widespread attention. The maximin constraint, a common notion of fairness adopted in the FIM problem, imposes a direct and intuitive requirement that asks the utility (influenced ratio within a group) of the worst-off group should be maximized. Although the objective of FIM under maximin constraint is conceptually straightforward, the development of efficient algorithms with strong theoretical guarantees remains an open challenge. The difficulty arises from the fact that the maximin objective does not satisfy submodularity, a key property for designing approximate algorithms in traditional influence maximization settings. In this paper, we address this challenge by proposing a two-step optimization framework consisting of Inner-group Maximization (IGM) and Across-group Maximization (AGM). We first prove that the influence spread within any individual group remains submodular, enabling effective optimization within groups. Based on this, IGM applies a greedy approach to pick high-quality seeds for each group. In the second step, AGM coordinates seed selection across groups by introducing two strategies: Uniform Selection (US) and Greedy Selection (GS). We prove that AGM-GS holds a $(1-1/e-\varepsilon)$ approximation to the optimal solution when groups are completely disconnected, while AGM-US guarantees a roughly $\frac{1}{m}(1-1/e-\varepsilon)$ lower bound regardless of the group structure, with $m$ denoting the number of groups.
Introduction
@@
Influence maximization (IM) (Li et al. 2018; Banerjee, Jenamani, and Pratihar 2020), a fundamental problem in social network analysis, aims to select a small set of seed nodes that maximizes the spread of influence in a network. While classical approaches primarily focus on maximizing total influence, recent work (Tsang et al. 2019; Fish et al. 2019; Farnadi, Babaki, and Gendreau 2020; Rui et al. 2023; Chowdhary et al. 2024; Wang et al. 2025) has addressed the issue of fairness in the IM problem, especially in scenarios where the population is divided into distinct demographic groups. In such a scenario, traditional IM methods may disproportionately benefit well-connected groups, leaving underrepresented groups with significantly lower influence coverage.
@@
To depict the fairness in the IM problem, existing studies have come up with various notions of fairness, including maximin fairness (Tsang et al. 2019; Fish et al. 2019), equity fairness (Farnadi, Babaki, and Gendreau 2020), welfare fairness (Rahmattalabi et al. 2021; Rui et al. 2023), concave fairness (Ali et al. 2023; Wang et al. 2025), etc. Among these notions, maximin fairness, also known as the maximin constraint, stands out for its simplicity and intuitive appeal. It imposes a clear requirement: the utility (i.e., the proportion of influenced individuals) of the worst-off group should be maximized. In other words, the algorithm should prioritize improving the influence spread for the most disadvantaged group, thereby ensuring a more balanced and fairer outcome.
Despite the simplicity of the maximin constraint, existing studies face challenges in designing efficient algorithms for Fair Influence Maximization (FIM) under this constraint. The fundamental difficulty lies in the fact that the maximin objective does not exhibit submodularity (Tsang et al. 2019), a crucial property that enables efficient approximation with provable guarantees in IM settings.
@@
To tackle the FIM problem under the maximin constraint, Tsang et al. (Tsang et al. 2019) applied a Frank-Wolfe style multiobjective optimization approach, treating the utility of each group as a separate objective. By setting a common utility threshold and gradually increasing it across all groups, their method aims to improve the worst-off group’s utility. However, their approach is computationally inefficient and was only evaluated on small networks with fewer than a hundred nodes in their original paper. Fish et al. (Fish et al. 2019) proposed a novel $\phi$-MEAN fairness, where setting $\phi=$-$\infty$ recovers the maximin objective. They introduced a Myopic approach, which iteratively selects the node with the lowest probability of being influenced. They also outlined a naive Greedy algorithm that, in each iteration, picks the node that maximizes the objective function as a seed. The naive Greedy achieves better objective values but at the cost of high computational complexity, whereas the Myopic approach performs better in terms of individual fairness, but is less effective for group-wise fairness. Rahmattalabi et al. (Rahmattalabi et al. 2019) reformulated the maximin objective into an approximate setting and addressed it using Mixed-Integer Linear Programming (MILP). While this approach is theoretically sound, it suffers from scalability issues and is limited in practice to networks with fewer than two hundred nodes. In summary, existing methods face significant challenges in designing efficient algorithms with provable approximation guarantees under the maximin constraint, hindering their practical applications to large-scale networks.
@@
In this paper, we address this challenge by proposing efficient approximation algorithms for FIM under the maximin constraint, along with provable theoretical guarantees. We first prove that although the global maximin objective is not submodular, the influence spread within each individual group still satisfies submodularity. Leveraging this property, we design a two-step optimization framework for seed selection: the first step performs inner-group maximization, and the second step coordinates across-group maximization to improve the utility of the worst-off group. In the second step, we develop two algorithms with different coordination strategies, each offering provable approximation guarantees under different group structures.
@@
  • AGM introduces two seed selection strategies: Uniform Selection (AGM-US) and Greedy Selection (AGM-GS). AGM-US coordinates seed selection in a column-wise manner, establishing an approximation guarantee of roughly $\frac{1}{m}(1-1/e-\varepsilon)$ to the optimal solution regardless of group structure, with $m$ denoting the number of groups. AGM-GS employs a greedy manner, providing an $(1-1/e-\varepsilon)$ approximation only when the groups are completely disconnected.
  • Experimental results on five real-world networks with different group structures demonstrate that our proposed framework consistently produces superior seed sets across all settings. In particular, AGM-GS typically outperforms AGM-US, achieving better empirical performance than the proven theoretical bound of AGM-US while incurring a lower price of fairness.
Preliminaries
A network can be modeled as a graph $G=(V,E)$, where $V$ represents the set of nodes and $E\subseteq V\times V$ denotes the set of edges that connect node pairs. $\mathcal{C}=\{c_{1},c_{2},\cdots,c_{m}\}$ denotes the non-overlapping group (community) structure in $G$ with $V_{i}$ and $E_{i}$ denote the set of nodes and edges in $c_{i}$, respectively. Thus, it holds $V_{i}\cap V_{j}=\varnothing$ for any $c_{i},c_{j}\in\mathcal{C},i\neq j$. An edge $e=(u,v)\in E$ with $u\in c_{i},v\in c_{j}$ is called an inner-group edge if $c_{i}=c_{j}$. On the contrary, the edge is called a cross-group edge if $c_{i}\neq c_{j}$. In this paper, we define group connectivity $\rho(G,\mathcal{C})$ as the ratio between the number of inner-group edges and the number of total edges. Formally, $\rho(G,\mathcal{C})=\frac{1}{|E|}\sum_{c\in\mathcal{C}}\sum_{u\in V_{c},v\notin V_{c}}\chi(u,v)$, where $\chi(u,v)=1$ if $(u,v)\in E$, and $\chi(u,v)=0$ otherwise.
Independent Cascade Model
The Independent Cascade (IC) model, originally proposed by Kempe et al. (Kempe, Kleinberg, and Tardos 2003), is one of the most widely adopted information diffusion models in the field of Influence Maximization (IM). The input of Ic model includes a graph $G=(V,E)$ and a probability matrix $P\in[0,1]^{|V|*|V|}$, where each edge $e=(u,v)\in E$ is associated with a probability $p(e)=p(u,v)\in P$, representing the likelihood that node $u$ activates node $v$. In this study, we adopt a common setting for $P$, where $p(u,v)=1/d^{-}(v)$ with $d^{-}(v)$ represents the in-degree of node $v$ (Chen, Wang, and Wang 2010; Jung, Heo, and Chen 2012; Tang, Shi, and Xiao 2015; Xie et al. 2024; Rui et al. 2025).
@@
Influence Spread
The influence spread of a node set $S$, usually denoted as $\sigma(S)$, is a set function that represents the expected number of nodes activated during the information diffusion process. Formally, it is defined as $\sigma(S)=\mathbb{E}[|A_{0}\cup A_{1}\cup\cdots\cup A_{t_{\text{ed}}}|]$. In addition, we define $\sigma_{c}(S)$ as the inner influence spread for group $c$, where only activated nodes in $c$ are counted, namely, $\sigma_{c}(S)=\mathbb{E}[|A_{0}\cup A_{1}\cup\cdots\cup A_{t_{\text{ed}}}\cap V_{c}|]$. The utility of group $c$ is then defined as $\boldsymbol{u}_{c}(S)=\sigma_{c}(S)/|V_{c}|$, representing the influenced ration within $c$.
Live-edge Graph
@@
Inner-group Maximization (IGM)
@@
For a graph $G=(V,E)$ and its associated probability matrix $P$, let $\mathcal{L}$ be a randomly sampled live-edge graph under the IC model and $\Gamma_{c}(S,\mathcal{L})$ denote the set of nodes in group $c$ that are reachable from $S$ within $\mathcal{L}$. Then, for any $S\subseteq T\subseteq V$, it holds $\Gamma_{c}(S,\mathcal{L})\subseteq\Gamma_{c}(T,\mathcal{L})$. Let $v\in V,v\notin T$ be a random node, then it meets
@@
Recall that $\sigma(S)=\sum_{\mathcal{L}\in\Omega(\mathcal{L},G)}\Pr\{\mathcal{L}|G\}\cdot|\Gamma(S,\mathcal{L})|$, we have $\sigma_{c}(S)=\sum_{\mathcal{L}\in\Omega(\mathcal{L},G)}\Pr\{\mathcal{L}|G\}\cdot|\Gamma_{c}(S,\mathcal{L})|$, which leads to
@@
which indicates $\sigma_{c}(\cdot):V\rightarrow\mathbb{R}$ is submodular, and thus concludes the proof. ∎
In addition, $\sigma_{c}$ is obviously non-negative and monotonic according to its definition. Therefore, by applying a greedy approach that iteratively selects the node bringing the maximum gain to $\sigma_{c}$ into the seed set $S$, it will hold $\sigma_{c}(S)\geq(1-1/e)\cdot OPT$ (Kempe, Kleinberg, and Tardos 2003) with $OPT$ denoting the optimal result.
Thus, our proposed Inner-group Maximization (IGM) applies IMM (Tang, Shi, and Xiao 2015) within each group $c\in\mathcal{C}$ to maximize their inner influence spread $\sigma_{c}(S^{c})$ and obtain its seed set, denoted as $S^{c}=\{s^{c}_{1},s^{c}_{2},\cdots,s^{c}_{k}\}$. Due to the fact that calculating $\sigma_{c}$ is #P-hard (Kempe, Kleinberg, and Tardos 2003), IMM estimates $\sigma_{c}$ based on the Reverse Influence Sampling (RIS) (Borgs et al. 2014; Tang, Xiao, and Shi 2014) approach. By generating $\theta_{c}=\frac{2n\cdot\left((1-1/e)\cdot\sqrt{\ell\ln|V_{c}|+\ln 4}+\sqrt{(1-1/e)\cdot\left(\ln\binom{|V_{c}|}{k}+\ell\ln{|V_{c}|}+\ln 4\right)}\right)^{2}}{OPT_{c}\cdot\varepsilon^{2}}$ Reverse Reachable (RR) sets (Chen 2018), the seed set $S^{c}$ identified via IMM guarantees the approximation of $\sigma_{c}(S^{c})\geq(1-1/e-\varepsilon)\cdot OPT_{c}$ with a probability at least $1-1/|V_{c}|^{\ell}$. Let $S^{*c}$ denote the optimal seed set which leads to $\sigma(S^{*c})=OPT_{c}$, it is essential to note that the seeds in $S^{*c}$ do not necessarily have to be inside $c$, which means it is a global solution. The same formulation also holds for $S^{c}$.
@@
Across-group Maimixzation (AGM)
@@
$\begin{split}\Phi(S)&=\boldsymbol{u}_{c^{\#}}(S)\\ &\geq(\frac{1}{m}-\xi)(1-1/e-\varepsilon)\boldsymbol{u}_{c^{\#}}(S^{*})\\ &\geq(\frac{1}{m}-\xi)(1-1/e-\varepsilon)\boldsymbol{u}_{c^{*}}(S^{*})\\ &\geq(\frac{1}{m}-\xi)(1-1/e-\varepsilon)\Phi(S^{*}).\end{split}$
@@
providing a more tight and empirical bound.
@@
Assume the number of seeds selected in each group $c$ from $S^{*}$ is $b_{c}$, i.e., $|S^{*}\cup V_{c}|=b_{c}$, and the seed sets output by IGM is $\mathcal{S}=\{S^{c_{1}},S^{c_{2}},\cdots,S^{c_{m}}\}$. Let $S^{c}_{r}=\{s^{c}_{1},s^{c}_{2},\cdots,s^{c}_{r}\}$ ($r\leq k$) denoting the first $r$ seeds in $S^{c}$ for each group $c$, $c\in\mathcal{C}$, then we have
$\begin{split}\Phi(S^{c}_{b_{c}})&\geq(1-1/e-\varepsilon)\Phi(S^{*c}_{b_{c}})\\ &=(1-1/e-\varepsilon)\Phi(S^{*}_{b_{c}}).\end{split}$ (4)
@@
$\begin{split}\Phi(S)&=\boldsymbol{u}_{c^{\#}}(S)\\ &=\boldsymbol{u}_{c^{\#}}(S^{c^{\#}}_{d_{c^{\#}}})\\ &\geq\boldsymbol{u}_{c^{\#}}(S^{c^{\#}}_{b_{c^{\#}}})\\ &\geq(1-1/e-\varepsilon)\boldsymbol{u}_{c^{\#}}(S^{*c^{\#}}_{b_{c^{\#}}})\\ &\geq(1-1/e-\varepsilon)\boldsymbol{u}_{c^{*}}(S^{*c^{*}}_{b_{c^{*}}})\\ &=(1-1/e-\varepsilon)\Phi(S^{*}).\end{split}$ (5)
@@
$\begin{split}\boldsymbol{u}_{c^{\prime}}(S^{c^{\prime}}_{d_{c^{\prime}}-1})&\geq\boldsymbol{u}_{c^{\prime}}(S^{c^{\prime}}_{b_{c^{\prime}}})\\ &\geq(1-1/e-\varepsilon)\boldsymbol{u}_{c^{\prime}}(S^{*c^{\prime}}_{b_{c^{\prime}}})\\ &\geq(1-1/e-\varepsilon)\boldsymbol{u}_{c^{*}}(S^{*c^{*}}_{b_{c^{*}}})\\ &=(1-1/e-\varepsilon)\Phi(S).\end{split}$ (6)
@@
$\begin{split}\Phi(S)&=\boldsymbol{u}_{c^{\#}}(S)\\ &=\boldsymbol{u}_{c^{\#}}(S^{c^{\#}}_{d_{c^{\#}}})\\ &\geq\boldsymbol{u}_{c^{\prime}}(S^{c^{\prime}}_{d_{c^{\prime}}-1})\\ &\geq(1-1/e-\varepsilon)\boldsymbol{u}_{c^{\prime}}(S^{*c^{\prime}}_{b_{c^{\prime}}})\\ &\geq(1-1/e-\varepsilon)\boldsymbol{u}_{c^{*}}(S^{*c^{*}}_{b_{c^{*}}})\\ &\geq(1-1/e-\varepsilon)\Phi(S^{*}).\end{split}$
@@
Therefore, $\Phi(S)\geq(1-1/e-\varepsilon)\Phi(S)$ always holds for both cases, which concludes the proof. ∎
Discussion on AGM and $\rho$

### v2

@@ Abstract
Fair Influence Maximization (FIM) seeks to mitigate disparities in influence across different groups and has recently garnered increasing attention. A widely adopted notion of fairness in FIM is the maximin constraint, which directly requires maximizing the utility (influenced ratio within a group) of the worst-off group. Despite its intuitive formulation, designing efficient algorithms with strong theoretical guarantees remains challenging, as the maximin objective does not satisfy submodularity, a key property for designing approximate algorithms in traditional influence maximization settings. In this paper, we address this challenge by proposing a two-step optimization framework consisting of Inner-group Maximization (IGM) and Across-group Maximization (AGM). We first prove that the influence spread within any individual group remains submodular, enabling effective optimization within groups. Based on this, IGM applies a greedy approach to pick high-quality seeds for each group. In the second step, AGM coordinates seed selection across groups by introducing two strategies: Uniform Selection (US) and Greedy Selection (GS). We prove that AGM-GS holds a $(1-1/e-\varepsilon)$ approximation to the optimal solution when groups are completely disconnected, while AGM-US guarantees a roughly $\frac{1}{m}(1-1/e-\varepsilon)$ lower bound regardless of the group structure, with $m$ denoting the number of groups.
1 Introduction
@@
Influence maximization (IM) (Li et al., 2018; Banerjee et al., 2020), a fundamental problem in social network analysis, aims to select a small set of seed nodes that maximizes the spread of influence in a network. While classical approaches primarily focus on maximizing total influence, recent work (Tsang et al., 2019; Fish et al., 2019; Farnadi et al., 2020; Rui et al., 2023; Chowdhary et al., 2024; Wang et al., 2025) has addressed the issue of fairness in the IM problem, especially in scenarios where the population is divided into distinct demographic groups. In such a scenario, traditional IM methods may disproportionately benefit well-connected groups, leaving underrepresented groups with significantly lower influence coverage.
@@
To depict the fairness in the IM problem, existing studies have come up with various notions of fairness, including maximin fairness (Tsang et al., 2019; Fish et al., 2019), equity fairness (Farnadi et al., 2020), welfare fairness (Rahmattalabi et al., 2021; Rui et al., 2023), concave fairness (Ali et al., 2023; Wang et al., 2025), etc. Among these notions, maximin fairness, also known as the maximin constraint, stands out for its simplicity and intuitive appeal. It imposes a clear requirement: the utility (i.e., the proportion of influenced individuals) of the worst-off group should be maximized. In other words, the algorithm should prioritize improving the influence spread for the most disadvantaged group, thereby ensuring a more balanced and fairer outcome.
@@
To tackle the FIM problem under the maximin constraint, Tsang et al. (2019) applied a Frank-Wolfe style multiobjective optimization approach, treating the utility of each group as a separate objective. However, their approach is computationally inefficient and was only evaluated on small networks with fewer than a hundred nodes in their original paper. Fish et al. (2019) proposed a novel $\phi$-MEAN fairness, where setting $\phi=$-$\infty$ recovers the maximin objective. They outlined a naive Greedy algorithm that, in each iteration, picks the node that maximizes the objective function as a seed. The naive Greedy achieves better objective values but at the cost of high computational complexity. Rahmattalabi et al. (2019) reformulated the maximin objective into an approximate setting and addressed it using Mixed-Integer Linear Programming (MILP). While the approach is theoretically sound, it suffers from scalability issues and is limited in practice to networks with fewer than two hundred nodes. Becker et al. (2022) relaxed maximin by allowing for randomized strategies, producing distributions over sets rather than a deterministic seed set, which differs from the study in this paper. Recently, Spaeh and Miyauchi (2025) revisited the maximin problem as a multiobjective submodular maximization task and proposed the LP-Greedy algorithm, which relies on the Gurobi Optimizer (Gurobi Optimization LLC., 2025). While LP-Greedy achieves a ($1-1/e-\varepsilon$)-approximation under strict conditions (such as when the optimal solution is known or the solution of LP exceeds the optimal one), it remains computationally inefficient on large-scale networks.
In summary, despite the simplicity of the maximin constraint, existing studies face challenges in designing efficient algorithms with provable approximation guarantees under this constraint, hindering their practical applications to large-scale networks. The fundamental difficulty lies in the fact that the maximin objective does not exhibit submodularity (Tsang et al., 2019), a crucial property that enables efficient approximation with provable guarantees in traditional IM settings.
@@
In this paper, we address this challenge by proposing efficient approximation algorithms for FIM under the maximin constraint, along with provable theoretical guarantees. We introduce a two-step optimization framework: the first step performs inner-group maximization, where we show that influence spread within each group remains submodular though the maximin objective does not; and (2) the second step coordinates across-group maximization to improve the utility of the worst-off group with different coordination strategies, each offering provable approximation guarantees.
@@
  • AGM introduces two seed selection strategies: Uniform Selection (AGM-US) and Greedy Selection (AGM-GS). AGM-US coordinates seed selection in a column-wise manner, establishing an approximation guarantee of roughly $\frac{1}{m}(1-1/e-\varepsilon)$ to the optimal solution regardless of group structure, with $m$ denoting the number of groups. AGM-GS employs a greedy manner, providing an $(1-1/e-\varepsilon)$ approximation when the groups are completely disconnected.
  • Experimental results on seven real-world networks with different group structures demonstrate that our proposed framework consistently produces superior seed sets across all settings. In particular, AGM-GS typically outperforms AGM-US, achieving better empirical performance than the proven theoretical bound of AGM-US while incurring a lower price of fairness.
2 Preliminaries
A network can be modeled as a graph $G=(V,E)$, where $V$ represents the set of nodes and $E\subseteq V\times V$ denotes the set of edges that connect node pairs. $\mathcal{C}=\{c_{1},c_{2},\cdots,c_{m}\}$ denotes the non-overlapping group (community) structure in $G$ with $V_{i}$ denote the set of nodes in $c_{i}$. Thus, it holds $V_{i}\cap V_{j}=\varnothing$ for any $c_{i}\neq c_{j}$. An edge $e=(u,v)\in E$ with $u\in c_{i},v\in c_{j}$ is called an inner-group edge if $c_{i}=c_{j}$. On the contrary, the edge is called a cross-group edge if $c_{i}\neq c_{j}$. In this paper, we define group connectivity $\rho(G,\mathcal{C})$ as the ratio between the number of cross-group edges and the number of total edges. Formally, $\rho(G,\mathcal{C})=\frac{1}{|E|}\sum_{c\in\mathcal{C}}\sum_{u\in V_{c},v\notin V_{c}}\chi(u,v)$, where $\chi(u,v)=1$ if $(u,v)\in E$, and $\chi(u,v)=0$ otherwise.
2.1 Independent Cascade Model
The Independent Cascade (IC) model, originally proposed by Kempe et al. (2003), is one of the most widely adopted information diffusion models in the field of Influence Maximization (IM). The input of IC model includes a graph $G=(V,E)$ and a probability matrix $P\in[0,1]^{|V|*|V|}$, where each edge $e=(u,v)\in E$ is associated with a probability $p(e)=p(u,v)\in P$, representing the likelihood that node $u$ activates node $v$. In this study, we adopt a common setting for $P$, where $p(u,v)=1/d^{-}(v)$ with $d^{-}(v)$ represents the in-degree of node $v$ (Chen et al., 2010; Jung et al., 2012; Tang et al., 2015; Xie et al., 2024; Rui et al., 2025).
@@ 2.1 Independent Cascade Model
Note that our work supports the triggering model (Kempe et al., 2003), a general model that includes the IC model as a special case. We adopt IC solely for ease of comparison.
2.2 Influence Spread
The influence spread of a node set $S$, usually denoted as $\sigma(S)$, is a set function that represents the expected number of nodes activated. Formally, it is defined as $\sigma(S)=\mathbb{E}[|A_{0}\cup A_{1}\cup\cdots\cup A_{t_{\text{ed}}}|]$. In addition, we define $\sigma_{c}(S)$ as the inner influence spread for group $c$, where only activated nodes in $c$ are counted, namely, $\sigma_{c}(S)=\mathbb{E}[|A_{0}\cup A_{1}\cup\cdots\cup A_{t_{\text{ed}}}\cap V_{c}|]$. The utility of group $c$ is then defined as $\boldsymbol{u}_{c}(S)=\sigma_{c}(S)/|V_{c}|$, representing the influenced ratio within $c$.
2.3 Live-edge Graph
@@ 3 Method
3.1 Inner-group Maximization (IGM)
@@
For a graph $G=(V,E)$ and its associated probability matrix $P$, let $\mathcal{L}$ be a randomly sampled live-edge graph under the IC model and $\Gamma_{c}(S,\mathcal{L})$ denote the set of nodes in group $c$ that are reachable from $S$ within $\mathcal{L}$. Then, for any $S\subseteq T\subseteq V$, it holds $\Gamma_{c}(S,\mathcal{L})\subseteq\Gamma_{c}(T,\mathcal{L})$. Let $v\in V,v\notin T$ be a random node, then it holds
@@
Recall that $\sigma(S)=\sum_{\mathcal{L}\in\Omega(\mathcal{L},G)}\Pr\{\mathcal{L}|G\}\cdot|\Gamma(S,\mathcal{L})|$, we have $\sigma_{c}(S)=\sum_{\mathcal{L}\in\Omega(\mathcal{L},G)}\Pr\{\mathcal{L}|G\}\cdot|\Gamma_{c}(S,\mathcal{L})|$, leading to
@@
which concludes the proof. ∎
In addition, $\sigma_{c}$ is obviously non-negative and monotonic according to its definition. Therefore, by applying a greedy approach that iteratively adds the node bringing the maximum gain w.r.t $\sigma_{c}$ to $S$, it will hold $\sigma_{c}(S)\geq(1-1/e)\cdot OPT$ (Kempe et al., 2003) with $OPT$ denoting the optimal result.
Thus, our proposed Inner-group Maximization (IGM) applies IMM (Tang et al., 2015) within each group $c\in\mathcal{C}$ to maximize their inner influence spread $\sigma_{c}(S^{c})$ and obtain its correspoding seed set, denoted as $S^{c}=\{s^{c}_{1},s^{c}_{2},\cdots,s^{c}_{k}\}$. Note that $s_{i}^{c}$ may lie outside group $c$. Due to the fact that calculating $\sigma_{c}$ is #P-hard (Kempe et al., 2003), IMM estimates $\sigma_{c}$ based on the Reverse Influence Sampling (RIS) (Borgs et al., 2014; Tang et al., 2014) approach. By generating $\theta_{c}=\frac{2n\cdot\left((1-1/e)\cdot\sqrt{\ell\ln|V_{c}|+\ln 4}+\sqrt{(1-1/e)\cdot\left(\ln\binom{|V_{c}|}{k}+\ell\ln{|V_{c}|}+\ln 4\right)}\right)^{2}}{OPT_{c}\cdot\varepsilon^{2}}$ Reverse Reachable (RR) sets (Chen, 2018), where $\varepsilon$ (usually 0.1) and $\ell$ (usually 1) are two manual parameters, the seed set $S^{c}$ identified via IMM guarantees the approximation of $\sigma_{c}(S^{c})\geq(1-1/e-\varepsilon)\cdot OPT_{c}$ with a probability at least $1-1/|V_{c}|^{\ell}$. Let $S^{*c}$ denote the optimal seed set which leads to $\sigma(S^{*c})=OPT_{c}$. It is essential to note that the seeds in $S^{*c}$ do not necessarily have to be inside $c$, which means it is a global solution. The same formulation also holds for $S^{c}$.
@@
3.2 Across-group Maximization (AGM)
@@ 3.2.1 Uniform Selection
$\begin{split}\Phi(S)=\boldsymbol{u}_{c^{\#}}(S)&\geq(\frac{1}{m}-\xi)(1-1/e-\varepsilon)\boldsymbol{u}_{c^{\#}}(S^{*})\\ &\geq(\frac{1}{m}-\xi)(1-1/e-\varepsilon)\boldsymbol{u}_{c^{*}}(S^{*})\\ &\geq(\frac{1}{m}-\xi)(1-1/e-\varepsilon)\Phi(S^{*}).\end{split}$
@@
providing a tighter and empirical bound.
3.2.2 Greedy Selection
@@
Assume the number of seeds selected in each group $c$ from $S^{*}$ is $b_{c}$, i.e., $|S^{*}\cap V_{c}|=b_{c}$, and the seed sets output by IGM is $\mathcal{S}=\{S^{c_{1}},S^{c_{2}},\cdots,S^{c_{m}}\}$. Let $S^{c}_{r}=\{s^{c}_{1},s^{c}_{2},\cdots,s^{c}_{r}\}$ ($r\leq k$) denoting the first $r$ seeds in $S^{c}$ for each group $c$, $c\in\mathcal{C}$, then we have
$\Phi(S^{c}_{b_{c}})\geq(1-1/e-\varepsilon)\Phi(S^{*c}_{b_{c}})=(1-1/e-\varepsilon)\Phi(S^{*}_{b_{c}}).$ (4)
@@
$\begin{split}\Phi(S)&=\boldsymbol{u}_{c^{\#}}(S)=\boldsymbol{u}_{c^{\#}}(S^{c^{\#}}_{d_{c^{\#}}})\\ &\geq\boldsymbol{u}_{c^{\#}}(S^{c^{\#}}_{b_{c^{\#}}})\geq(1-1/e-\varepsilon)\boldsymbol{u}_{c^{\#}}(S^{*c^{\#}}_{b_{c^{\#}}})\\ &\geq(1-1/e-\varepsilon)\boldsymbol{u}_{c^{*}}(S^{*c^{*}}_{b_{c^{*}}})\\ &=(1-1/e-\varepsilon)\Phi(S^{*}).\end{split}$
@@
$\begin{split}\boldsymbol{u}_{c^{\prime}}(S^{c^{\prime}}_{d_{c^{\prime}}-1})&\geq\boldsymbol{u}_{c^{\prime}}(S^{c^{\prime}}_{b_{c^{\prime}}})\\ &\geq(1-1/e-\varepsilon)\boldsymbol{u}_{c^{\prime}}(S^{*c^{\prime}}_{b_{c^{\prime}}})\\ &\geq(1-1/e-\varepsilon)\boldsymbol{u}_{c^{*}}(S^{*c^{*}}_{b_{c^{*}}})\\ &=(1-1/e-\varepsilon)\Phi(S).\end{split}$
@@
$\begin{split}\Phi(S)=\boldsymbol{u}_{c^{\#}}(S)&=\boldsymbol{u}_{c^{\#}}(S^{c^{\#}}_{d_{c^{\#}}})\\ &\geq\boldsymbol{u}_{c^{\prime}}(S^{c^{\prime}}_{d_{c^{\prime}}-1})\\ &\geq(1-1/e-\varepsilon)\boldsymbol{u}_{c^{\prime}}(S^{*c^{\prime}}_{b_{c^{\prime}}})\\ &\geq(1-1/e-\varepsilon)\boldsymbol{u}_{c^{*}}(S^{*c^{*}}_{b_{c^{*}}})\\ &\geq(1-1/e-\varepsilon)\Phi(S^{*}).\end{split}$
@@
Therefore, $\Phi(S)\geq(1-1/e-\varepsilon)\Phi(S^{*})$ always holds for both cases, which concludes the proof. ∎
It should be noted that our proposed framework is most effective when the number of communities is small, which is our intended setting and also where fairness (e.g., gender or ethnicity) is most meaningful. Besides, it also requires $k>m$, as this avoids the regime where the problem becomes inapproximable (Krause et al., 2008).
3.3 Discussion on AGM and $\rho$